# Integrating Conductor Health into Dynamic Line Rating and Unit Commitment under Uncertainty

### v2

@@ IV-B Case study I - Value of Conductor Health Awareness
((a))
((b))
@@
IV-B1 Overall Performance Comparison
Table I presents a breakdown of the total cost of three models: baseline SLR, baseline DLR and CHA-UC. The units are normalized to one day. The day-ahead cost corresponds to the objective value of (6), which includes UC costs, fuel costs and the day-ahead load shedding costs.
CHA-UC achieves the lowest annual total cost of 21.57 M\$, reducing costs by 0.8% and 2.2% compared to the SLR and DLR baselines, respectively. This improvement reflects its ability to balance the risk of ETO against reduced operation costs. Although CHA-UC adopts DLR, it applies a conservative line flow than the baseline by accounting for expected conductor depreciation cost. As a result, the DLR baseline yields the lowest day-ahead cost, followed by CHA-UC and then SLR. WP curtailments are reduced by 88% (9.1 v.s. 77.1) when DLR is used instead of SLR, as it relieves grid congestion. CHA-UC retains part of this advantage, though its robustness discards some of it, leading to a 83% reduction (12.7 v.s. 77.1). On the other hand, depreciation costs were lowest for CHA-UC followed by SLR and DLR indicating that the CHA-UC preserves conductor lifespan most effectively. Finally, reserves are only activated in CHA-UC, as the SLR and DLR baselines involve no uncertainty in this case study.
IV-B2 Risk of ETO
@@
Figure 6 visualizes the annual line temperature from the three models. In spring, fall and winter, the line temperatures were high in order of DLR baseline, CHA-UC and SLR baseline. This order originates from DLR not considering stochasticity with high ratings, CHA-UC considering stochasticity with high ratings, and SLR using a uniform low rating. However, during summer, the overall line temperatures increase and the temperature of SLR exceeds CHA-UC. This is because the hot operating conditions of summer increases the risk of SLR. Since summer weathers are close to the conservative conditions for SLR calculation, the “conservative” SLR becomes risky as well. Thus, CHA-UC chooses a more conservative loading, shown in the day-ahead cost of 29.96 M\$, which is higher than 29.92 M\$ of SLR baseline. A zoom-in of the scatter plot exhibits the results during this period. The temperatures from SLR was stamped more commonly over the thermal limit compared to CHA-UC, although CHA-UC has some outliers. This resulted in a CHA-UC depreciation cost of 0.26 M\$ which was 3.3 times and 4.8 times less than SLR and DLR baseline, respectively.
IV-B3 Effect of CHA-UC to Commitment Decisions
@@
The annual average of commitment decisions made by DLR baseline is illustrated on a Texan map in Fig. 7(a). The cheapest units which comply with UC and DC PF constraints are committed, leading to aggressive line loading. The employment of CHA-UC leads to committing additional costly units to hedge the risks of ETO, leading to an overall relief of line loading in DLR lines, as presented in Fig. 7(b). For example, line 74, a key line connecting the south-west region and Dallas, is highly loaded in a direction to Dallas (Bus 22 to Bus 94) with an annual average of 83.5% (normalized by line capacity) when operated by DLR baseline. The transition to CHA-UC led to a 8.2%p loading relief. This contributes to an annual conductor health relief of 16.2%, which translates to a 13.3 M\$ reduction in depreciation costs. Here, the commitment decisions took a major role; the average commitment on bus 94 increased by 22%, and the commitment on bus 30, adjacent to the east of bus 94, increased by 28%, leading to the reduction of power flow to Dallas and the stress on line 74.
IV-C Case study II - Impact of Correlated Forecast Errors
@@
Case study 1 has shown the value of conductor health awareness. Case study 2 involves adding WP forecast errors. To handle wind generation forecast errors, a minimum reserve constraint of 5 GW (95% VaR of 2022 WP forecast error) is enforced.
((a))
((b))
((c))
@@
Figure 8: (Case Study II) Lineflow, temperature and depreciation cost of line 74 at day 60 from baseline SLR, DLR and CHA-UC. (a) the day-ahead lineflow. Day-ahead DLR values and the security margin are colored orange. (b) real-time lineflow, obtained as a result of the operation stage. (c) The resultant conductor line temperature in lines and depreciation costs in bars, obtained through post-hoc analysis.
IV-C1 Overall Performance Comparison
@@
The costs of the annual simulation are loaded on the right-hand side of Table I. Overall, the trend of the three models from case study I was preserved. CHA-UC recorded the lowest annual total cost, 25.5 M\$, contributing to a 25.2% and 19.8% reduction compared to DLR and SLR, respectively. However, unlike case study I, CHA-UC always incurred the highest day-ahead cost and it recorded the least reserve activation costs. This occurs as only CHA-UC accounts the stochasticity in WP forecasts in the day-ahead decision making.
IV-C2 Impact of WP forecast error
@@
In hour 22, the actual DLR values were much lower than the forecast, causing ETO for DLR baseline. On the other hand, in hour 23, the RT line flow of DLR baseline surpassed the security margin of the DA DLR forecast, due to generation adjustments to balance WP forecast errors. In hour 23, WP and DLR forecast errors combined unfavorably to amplify the depreciation. The line temperature and depreciation cost due to RT line flow in Fig. 8(c) visualize the impact of DLR and WP forecast errors. Counterintuitively, the temperature and depreciation costs are reversed in the two hours. This happens due to the diminishing depreciation cost, displayed in Fig. 3; the conductor loses sensitivity to ETO after high-temperature annealing.
IV-C3 Impact of Correlated WP and DLR Forecast Error
Five out of total six inter-asset correlations appeared from adjacent lines and wind farms. Two pairs of adjacent wind farms and lines were selected for deeper analysis. Table II lists the pairs; the two pairs have same signs of 1) correlation coefficients and 2) generation shift factor (GSF) of the WP bus to the line, but have different signs of average lineflow.
TABLE II: (Case Study II) Pairs of correlated wind generators and lines studied. The correlation was captured using Spearman’s correlation coefficients and the average line flow using DLR baseline is given. GSF was obtained by setting the paired bus as the from-bus.
Pair Wind Farm Line Correlation GSF Avg. Lineflow [pu]
1 8 74 0.38 0.26 -10.3
2 35 252 0.35 0.32 9.0
((a))
((b))
@@ IV-C3 Impact of Correlated WP and DLR Forecast Error
Figure 9 illustrates the heatmap of line temperature under baseline DLR. Forecast errors are calculated by subtracting actual values from the forecast value, meaning that a larger forecast error represents a larger overestimation of the value. Since the GSF is positive in pair 1, the shortage of WP leads to a reduction of line flow. The negative line flow becomes smaller, leading positive WP forecast error (shortage) to stress the conductor thermal limit. WP forecast errors are often hedged by the DLR forecast errors. The underestimation of DLR leads to an offset, where the WP shortage in the fourth quadrant is covered by the abundant line capacity. The first quadrant becomes risk-intensive due to the aggravation of DLR overestimation. Conversely, for pair 2, negative WP forecast errors (abundancy) stress the conductor thermal limit, leading the second quadrant to be risk-intensive.
@@
An experiment regarding correlation proves that the CHA-UC is successfully reflecting the emergent effect of correlation. The scenario indices of the pairs were scrambled, changing the marginal minimally, so that the correlation of the scenarios becomes close to 0. This will lead to the inconsideration of correlation. Separate experiments done with two pairs, revealed in Table III, show that the original (correlation-considering) CHA-UC has kept a less conservative line flow for the risk-prone pair 1 and a more conservative line flow for the risk-hedging pair 2. This asserts that the CHA-UC successfully handles the effect of correlation.
TABLE III: (Case Study II) Annual average CHA-UC day-ahead lineflows before and after removing correlation from the scenario set
Pair CHA-UC [pu] CHA-UC w/o Corr. [pu] Conservativeness
1 -4.96 -5.23 Decreased
2 7.96 6.70 Increased
V Conclusion
This paper presents CHA-UC, a stochastic unit commitment problem that considers line depreciation from ETO, effective for DLR operation. The CHA-UC optimization contains a conductor temperature estimation module and a conductor depreciation cost function to co-optimize the operation cost and depreciation cost. In result, case studies showcase a 0.8% and 83% reduction of total costs and WP curtailments compared to SLR operation, respectively. DLR operation without risk consideration showed highest costs due to the large conductor depreciation costs. The soundness of CHA-UC was proven by comparing annual line temperatures with the spatial commitment differences. A correlation analysis of DLR and WP forecast errors reveals the emergent properties from interactions of correlation, GSF and line flow. CHA-UC handles this properly. The statistics and analysis provided in the case studies will be important resources for DLR integration. The computational simplicity of the conductor temperature linear approximation will pay off. Despite some assumptions, our model proves to be most effective in the day-ahead operation with DLR, and provides meaningful materials for future grid research.

### v4

@@ 4.2 Case study I - Value of Conductor Health Awareness
4.2.1 Overall Performance Comparison
Table 1 presents a breakdown of the total cost of four models: baseline SLR, baseline DLR, QRF and CHA-UC. The units are normalized to one day. The day-ahead cost corresponds to the objective value of (9), which includes UC costs, fuel costs and the day-ahead load shedding costs. The total cost is the sum of day-ahead costs, real-time costs and conductor depreciation costs. The benchmark risk-averse measure, QRF-UC does not show advantage over baseline SLR, since its lower DA costs (21.29 vs. 21.44) are offset by the higher depreciation cost (0.31 vs. 0.16). Simulations with higher confidence level (95%) shows even higher total costs. CHA-UC achieves the lowest annual total cost of 21.44 M\$/day, reducing costs by 2.05%, 0.74% and 0.74% compared to DLR, SLR and QRF, respectively. This improvement reflects its ability to balance the risk of ETO against reduced operation costs. Although CHA-UC adopts DLR, it applies a conservative line flow than the baseline by accounting for expected conductor depreciation cost. As a result, the DLR baseline yields the lowest day-ahead cost, followed by CHA-UC and then SLR. WP curtailments are reduced by 89.1% (9.1 v.s. 83.4) when DLR is used instead of SLR, as it relieves grid congestion. CHA-UC retains most of this advantage, though its robustness discards some of it, leading to a 85.0% reduction (12.5 v.s. 83.4). Using QRF retains the least of this advantage, leading to a 66.1% reduction (28.2 vs. 83.4). On the other hand, depreciation costs were lowest for CHA-UC followed by SLR and DLR indicating that the CHA-UC preserves conductor lifespan most effectively. Finally, reserves are only activated in CHA-UC, as the SLR and DLR baselines involve no uncertainty in this case study.
@@ 4.2.1 Overall Performance Comparison
4.2.2 Risk of ETO
@@
Figure 6(a) visualizes the annual line temperature from the three models. In spring, fall and winter, the line temperatures were high in order of DLR baseline, CHA-UC and SLR baseline. This order originates from DLR not considering stochasticity with high ratings, CHA-UC considering stochasticity with high ratings, and SLR using a uniform low rating. However, during summer, the overall line temperatures increase and the temperature of SLR exceeds CHA-UC. This is because the hot operating conditions of summer increases the risk of SLR. Since summer weathers are close to the conservative conditions for SLR calculation, the “conservative” SLR becomes risky as well. Thus, CHA-UC chooses a more conservative loading, shown in the day-ahead cost of 29.99 M\$/day, which is higher than 29.96 M\$/day of SLR baseline. A zoom-in of the scatter plot exhibits the results during this period. The temperatures from SLR was stamped more commonly over the thermal limit compared to CHA-UC, although CHA-UC has some outliers. This resulted in a CHA-UC depreciation cost of 0.13 M\$/day which was 3.5 times and 7.3 times less than SLR and DLR baseline, respectively.
On the other hand, QRF chooses an even more conservative operation in summer, costing 30.35 M\$/day, reflected in Fig. 6(b). However, the cost measures were relatively ineffective, leading to a higher depreciation cost compared to CHA-UC (0.19 vs. 0.13). Despite the moving average temperature of QRF being similar with CHA-UC, the QRF outliers tend to show more common high temperature ETOs. The distribution in the zoom-in area shows that CHA-UC exploits the knowledge on the state of the conductor, balancing low-temperature ETOs with reduced operation costs, while QRF-UC frequently operates in higher temperature regions. This results in higher conductor depreciation costs of QRF albeit the higher expenditure for conservativeness.
4.2.3 Effect of CHA-UC to Commitment Decisions
@@
The annual average of commitment decisions made by DLR baseline is illustrated on a Texan map in Fig. 7(a). The cheapest units which comply with UC and DC PF constraints are committed, leading to aggressive line loading. The employment of CHA-UC leads to committing additional costly units to hedge the risks of ETO, leading to an overall relief of line loading in DLR lines, as presented in Fig. 7(b). For example, line 74, a key line connecting the south-west region and Dallas, is highly loaded in a direction to Dallas (Bus 22 to Bus 94) with an annual average of 84.5% (normalized by line capacity) when operated by DLR baseline. The transition to CHA-UC led to an 8.7%p loading relief. This contributes to an annual conductor health relief of 14.0%, which translates to a 112.4 M\$ reduction in depreciation costs. Here, the commitment decisions took a major role; the annual average commitment on bus 94 increased by 39%, and the commitment on bus 30, adjacent to the east of bus 94, increased by 34%, leading to the reduction of power flow to Dallas and the stress on line 74.
4.3 Case study II - Impact of Correlated Forecast Errors
@@
Case study I has shown the value of conductor health awareness. Case study II involves adding WP forecast errors to see the effects of correlated forecast errors. To handle wind generation forecast errors, a minimum reserve constraint of 5 GW (95% VaR of 2022 WP forecast error) is enforced.
@@
Figure 8: (Case Study II) Lineflow, temperature and depreciation cost of line 74 at day 107 from baseline SLR, DLR and CHA-UC. (a) the day-ahead lineflow. Day-ahead DLR values and the security margin are colored orange. (b) real-time lineflow, obtained as a result of the operation stage. (c) The resultant conductor line temperature in lines and depreciation costs in bars, obtained through post-hoc analysis.
4.3.1 Overall Performance Comparison
@@
The costs of the annual simulation are loaded on the right-hand side of Table 1. Overall, the trend of the three models from case study I was preserved. CHA-UC recorded the lowest annual total cost, 25.3 M\$, contributing to a 34.2%, 25.5% and 26.1% reduction compared to DLR, SLR and QRF, respectively. However, unlike case study I, CHA-UC always incurred the highest day-ahead cost and it recorded the least reserve activation costs. This occurs because only CHA-UC accounts the stochasticity in WP forecasts in the day-ahead decision making.
4.3.2 Impact of WP forecast error
@@
In hour 13 and 19, the RT line flow of DLR baseline surpassed the security margin of the DA DLR forecast, due to generation adjustments to balance WP forecast errors. On the other hand, in hour 21, the actual DLR values were lower than the forecast margin, causing ETO for DLR baseline. In all ETO cases above, WP and DLR forecast errors combined unfavorably to amplify the depreciation. The line temperature and depreciation cost due to RT line flow in Fig. 8(c) visualize the impact of DLR and WP forecast errors. The depreciation cost incurred from hours 13 and 19 were 0.23 M\$ and 0.12 M\$ caused by ETOs of 98.1°C and 97.6°C, respectively. The relatively large difference in depreciation cost from a modest temperature difference comes from two ways: 1) the depreciation cost increases exponentially by temperature and 2) the diminishing depreciation cost shown in Fig. 2 reduces the sensitivity to ETO for hour 19 after the ETO in hour 13.
4.3.3 Impact of Correlated WP and DLR Forecast Error
Two pairs of adjacent wind farms and lines were selected for an analysis on inter-asset correlation. Table 2 lists the pairs; the two pairs have same signs of 1) correlation coefficients and 2) generation shift factor (GSF) of the WP bus to the line, but have different signs of average lineflow.
Table 2: (Case Study II) Pairs of correlated wind generators and lines studied. The correlation was captured using Spearman’s correlation coefficients and the average line flow using DLR baseline is given. GSF and line flow were obtained by setting the paired bus as the from-bus.
Pair Wind Farm Line Corr. GSF Avg. Lineflow [pu]
1 8 74 0.38 0.26 -12.2
2 35 252 0.35 0.32 8.0
@@ 4.3.3 Impact of Correlated WP and DLR Forecast Error
Figure 9 illustrates the heatmap of line temperature under baseline DLR. Forecast errors are calculated by subtracting actual values from the forecast value, meaning that a larger forecast error represents a larger overestimation of the value. Since the GSF is positive in pair 1, the shortage of WP leads to a reduction of line flow. The negative line flow becomes smaller, leading positive WP forecast error (shortage) to stress the conductor thermal limit.
WP forecast errors are often hedged by the DLR forecast errors. The underestimation of DLR leads to an offset, where the WP shortage in the fourth quadrant is covered by the abundant line capacity. The first quadrant becomes risk-intensive due to the aggravation of DLR overestimation. Conversely, for pair 2, negative WP forecast errors (abundancy) stress the conductor thermal limit, leading the second quadrant to be risk-intensive.
Table 3: (Case Study II) Annual average CHA-UC day-ahead lineflows before and after removing correlation from the scenario set
Pair CHA-UC [pu] CHA-UC w/o Corr. [pu] Conservativeness
1 -5.09 -5.38 Decreased
2 8.06 6.76 Increased
@@
An experiment regarding correlation proves that the CHA-UC is successfully reflecting the emergent effect of correlation. The scenario indices of the pairs were scrambled, changing the marginal minimally, so that the correlation of the scenarios becomes close to 0. This will lead to the inconsideration of correlation. Separate experiments done with two pairs, revealed in Table 3, show that the original (correlation-considering) CHA-UC has kept a less conservative line flow for the risk-prone pair 1 and a more conservative line flow for the risk-hedging pair 2. This asserts that the CHA-UC successfully handles the effect of correlation.
5 Conclusion
This paper presents CHA-UC, a stochastic unit commitment problem that considers line depreciation from ETO, effective for DLR operation. The CHA-UC optimization contains a conductor temperature estimation module and a risk-base conductor depreciation cost function to co-optimize the operation cost and depreciation cost. The cost of conductor depreciation is quantified through risk-base valuation. In result, case studies showcase a 0.74% and 85% reduction of total costs and WP curtailments compared to SLR operation, respectively. Moreover, it outperformed the baseline risk-averse measure, QRF-UC with a 0.74% total cost difference, while DLR operation without risk consideration showed highest costs due to the large conductor depreciation costs. The soundness of CHA-UC was proven by comparing annual line temperatures with the spatial commitment differences. A correlation analysis of DLR and WP forecast errors reveals the emergent properties from interactions of correlation, GSF and line flow. CHA-UC handles this properly. The statistics and analysis provided in the case studies will be important resources for DLR integration. The computational simplicity of the conductor temperature linear approximation will pay off. Despite some assumptions, our model proves to be most effective in the day-ahead operation with DLR, and provides meaningful materials for future grid research.
[…]
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.
@@ Acknowledgement
This work was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Korean Government (MSIT) (No. RS-2023-00210018) and KENTECH Research Grant (202500011A).